OpenStudy (anonymous):

Trig question - help please. Solve for all real numbers. (Equation in comments.)

OpenStudy (anonymous):

\[2\cos(x)=-\sqrt{3}\]

OpenStudy (e.mccormick):

OK... where are you having trouble with this?

OpenStudy (anonymous):

I solved and got 5pi/6. I'm not sure if I would add 2pi to get the rest of the answers, in which case would the answer be 5pi/6 + 2npi?

OpenStudy (e.mccormick):

Just \(\frac{5\pi}{6}\)?

OpenStudy (anonymous):

yup

OpenStudy (e.mccormick):